### v6

@@ 4.2 Object detection experiments on COCO2017
+ CBAMConv(r) 4.3 1.7 41.9 3
+ CAConv(r) 4.3 1.7 42.4 3
+ RFAConv(r) 4.5 1.8 43.3 3
YOLOv5s 15.9 7.1 48.9 3
+ CAMConv(r) 16 7.1 48.5 3.5
+ CBAMConv(r) 16 7.1 49 3.7
@@
YOLOv7-tiny 13.2 6.1 50.2 5
+ CAMConv(r) 13.2 6.1 50.3 5.4
+ CBAMConv(r) 13.2 6.1 50.1 5.4
+ CAConv(r) 13.2 6.1 50.5 5.4
+ RFAConv(r) 13.6 6.1 50.6 7.5
YOLOv8n 8.1 3 53.5 3
+ CAMConv(r) 8.1 3 52.8 3.1
+ CBAMConv(r) 8.2 3 53.3 3.1
+ CAConv(r) 8.2 3 53.8 2.9
+ RFAConv(r) 8.4 3.1 54 3.2
Table 6: Object detection mAP50 and mAP on the VOC7+12 validation set.
4.3 Object detection experiments on VOC7+12
Following most of the work, we also report mAP. As shown in Table 6. As with the previously obtained conclusions, in all experiments, after we replace some of the convolution operations in the network using RFAConv, the network gained significant improvement by adding only a small number of parameters and computational overhead. Meanwhile, RFA obtains an outstanding performance compared to other attentions. Moreover, in some experiments, in some experiments, we similarly experiment with networks constructed by RFCBAMConv and RFCAConv. The results again validate their advantages. Also compared with CBAMConv and CAConv, they still obtained better results.
@@ 4.4 Semantic segmentation experiments on VOC2012
In order to validate the advantages of our method again, we conduct semantic segmentation experiments on the VOC2012 dataset, selecting DeepLabplusV3 [48] and the backbone network Resnet18 to conduct related experiments. The pre-training weights of each backbone network are obtained in the ImageNet-1k experiment. We report the results for outputs at two different step sizes, 8 and 16, respectively. In the experiment, we found that the semantic segmentation network constructed by RFAConv achieved better results than the original model, but compared to CAConv, CAMConv, the performance of RFAConv is not good. After thinking, we assert that RFAConv lacks consideration for long-distance information, while semantic segmentation tasks rely on long-distance information. CAConv, CAMConv and CBAMConv capture long range information by global averaging pooling to obtain global information. Although CBAMConv produce poor results for semantic segmentation, the improved RFCBAM obtain the good performance. This again demonstrates that spatial attention can again improve network performance through our approach, which simply requires making spatial attention attend to the receptive-field spatial feature.
@@ 4.5 Discussions
As all the experiments have shown, RFAConv serves as an alternative to standard convolution, bringing significant improvements to classification, target detection, and semantic segmentation visual tasks at the cost of a small increase in parameters and computational effort. As shown in the classification visualization in Fig. 9 and Fig. 10, networks built based on RFAConv are able to better focus on important information and features. This is due to the fact that RFA takes into account the receptive-field spatial features and is able to highlight the importance of each feature within the receptive-field, while combining it with convolution to turn it into a non-parametric shared convolution operation. Also as mentioned earlier, increasing the focus of CA and CBAM to the receptive-field spatial feature can again improve performance. Thus RFCBAM,RFCA was designed by us to significantly improve the network performance. What is clear is that RFAConv is advantageous but not outstanding in semantic segmentation tasks. We think it is that the design of RFA does not take long range information into account. As Hou et al.[18] argued, semantic segmentation relies on long range information. In contrast to RFA, RFCBAM and RFCA take into account the receptive-field spatial features and consider long range information through maximum pooling and average pooling. So they obtain outstanding performance in every visual task. By carefully analyzing RFAConv, RFCBAMConv and RFCAConv, our task can put some existing spatial attention into the receptive field to enhance its performance.
@@ 5 Conclusion
By analyzing the standard convolution and spatial attention, we conclude that spatial attention mechanisms address the problem of parameter sharing and have the advantage of considering long-distance information. However, the performance of the spatial attention mechanism is limited for the large-size convolution kernel. To address this issue, we propose a novel attention mechanism called RFA and devise a novel convolution operation, which further improves network performance. We also emphasize the importance of directing attention to the receptive-field spatial feature to enhance network performance. Through extensive experiments, we demonstrate the effectiveness and advanced nature of our approach. Going forward, we hope that more and more spatial attention mechanisms will adopt our proposed approach to further improve performance, and we also hope that the novel convolution method will be widely adopted to enhance network performance.

### v7

@@ 4.4 Object detection experiments on VOC7+12 and Roboflow-100
+ CBAMConv(r) 4.3 1.7 41.9 3.0
+ CAConv(r) 4.3 1.7 42.4 3.0
+ TAConv(r) 4.5 1.8 42.5 2.8
+ MCAConv(r) 4.2 1.8 41.1 2.9
+ HAMConv(r) 4.3 1.8 41.3 2.9
+ RFAConv(r) 4.5 1.8 43.3 3.0
YOLOv5s 15.9 7.1 48.9 3.0
+ CAMConv(r) 16.0 7.1 48.5 3.5
+ CBAMConv(r) 16.0 7.1 49.0 3.7
@@
YOLOv7-tiny 13.2 6.1 50.2 2.5
+ CAMConv(r) 13.2 6.1 50.3 2.7
+ CBAMConv(r) 13.2 6.1 50.1 2.7
+ CAConv(r) 13.2 6.1 50.5 2.7
+ TAConv(r) 13.4 6.1 50.4 2.8
+ MCAConv(r) 13.2 6.1 50.2 2.8
+ HAMConv(r) 13.2 6.1 50.1 2.7
+ RFAConv(r) 13.6 6.1 50.6 3.8
YOLOv8n 8.1 3.0 53.5 3.0
+ CAMConv(r) 8.1 3.0 52.8 3.1
+ CBAMConv(r) 8.2 3.0 53.3 3.1
+ CAConv(r) 8.2 3.0 53.8 2.9
+ TAConv(r) 8.4 3.0 54.0 3.3
+ MCAConv(r) 8.1 3.0 52.6 3.1
+ RFAConv(r) 8.4 3.0 54.0 3.2
Following most of the work, we also report mAP. As shown in Table 8. As with the previously obtained conclusions, in all experiments, after we replace some of the convolution operations in the network using RFAConv, the network gained significant improvement by adding only a small number of parameters and computational overhead. Meanwhile, RFA obtains an outstanding performance compared to other attentions. Moreover, in some experiments, we similarly experiment with networks constructed by RFCBAMConv and RFCAConv. The results again validate their advantages. For a fair comparison of other attention work, TAConv, HAMConv and MCAConv are still compared in advanced detection networks. All the experiments demonstrate that our proposed method obtains better performance.
Compared to the VOC7+12 dataset, Roboflow-100 includes 100 datasets from different domains. Therefore, to further demonstrate the effectiveness and generalizability of the proposed method, we conduct corresponding experiments on the large-scale Roboflow-100 dataset. We select YOLOv5n as the baseline network, training each network for 100 epochs, and the final reported results are obtained by averaging the experimental results from the 100 datasets. The experimental results are shown in Table 9. Compared to other methods, our proposed method achieves the best detection accuracy by replacing the convolutions in YOLOv5. The object detection experiments on VOC7+12 and Roboflow-100 further demonstrate the good generalization of our method.
Table 9: Object detection experiment results based on the Roboflow-100 dataset.
Models FLOPS (G) Param (M) Precision (%) Recall (%) mAP50 (%) mAP (%)
YOLOv5n 4.2 1.7 64.85 59.09 56.52 34.28
+ CAMConv(r) 4.2 1.7 64.01 59.89 56.53 34.44
+ CBAMConv(r) 4.3 1.7 64.55 58.92 56.4 34.6
+ CAConv(r) 4.3 1.7 65.01 59.5 57.01 34.81
+ TAConv(r) 4.5 1.8 63.99 58.25 56.23 33.99
+ MCAConv(r) 4.2 1.8 64.8 59.6 56.57 34.35
+ HAMConv(r) 4.3 1.8 65.01 59.93 56.84 34.67
+ RFAConv(r) 4.5 1.8 65.32 60.15 57.24 35.16
+ RFCBAMConv(r) 4.5 1.8 65.52 60.42 57.91 35.56
+ RFCAonv(r) 4.6 1.8 65.44 60.31 57.85 35.37
@@ 4.5 Semantic segmentation experiments on VOC2012
In order to validate the advantages of our method again, we conduct semantic segmentation experiments on the VOC2012 dataset, selecting DeepLabplusV3 [39] and the backbone network ResNet18 to conduct related experiments. Since the pre-trained weights of the backbone network are not available for other attention works, we only train a few networks that are able to take the pre-trained weights of the backbone network in the ImageNet-1k experiments. We report the results for outputs at two different step sizes, 8 and 16, respectively.
@@
As shown in Table 10, we found that the semantic segmentation network constructed by RFAConv achieved better results than the original model, but compared to CAConv, CAMConv, the performance of RFAConv is not good. After thinking, we assert that RFAConv lacks consideration for long-distance information, while semantic segmentation tasks rely on long-distance information. CAConv, CAMConv and CBAMConv capture long range information by global averaging pooling to obtain global information. Although CBAMConv produce poor results for semantic segmentation, the improved RFCBAM obtain the good performance. This again demonstrates that spatial attention can again improve network performance through our approach, which simply requires making spatial attention attend to the receptive-field spatial feature. To further underscore the merits of our proposed approach, we randomly select 5 images from the VOC2012 dataset and input them into the ResNet18 model constructed with various convolution operations for segmentation visualization and analysis. As illustrated in Figure 12, it is evident that the networks built using RFCBAMConv and RFCAConv, excel in delineating the intricate contours of the target compared to other methods. As previously noted, the significance of the receptive-field spatial features is paramount, as highlighted in the segmentation visualization contrast between CAConv and RFCAConv. The distinguishing factor of RFCAConv, in contrast to CAConv, lies in its emphasis on leveraging the receptive-field spatial features to tackle convolution parameter sharing challenges. Consequently, RFCAConv demonstrates superior segmentation outcomes when compared to CAConv.
Figure 12: Segmentation visualisation results of ResNet18 constructed with different convolutions. It is evident that RFAConv and RFCBAMConv allow the network to segment the detailed contours of the target more clearly than the other methods. CAConv differs from RFCAConv in that the latter focuses on the receptive-field spatial features, which proves the importance of the receptive-field spatial features.
@@ 4.6 Discussions
In classification experiments, RFAConv demonstrates substantial accuracy improvements over baseline models employing standard convolutions, despite marginal increases in parameters and computational overhead, as quantified in Table 3. This performance gap stems from fundamental architectural differences: while standard convolutions apply static kernel parameters across all spatial positions, RFAConv implements location-adaptive parameter modulation through learned attention weights for each receptive-field slide. This dynamic mechanism enables superior adaptation to spatial variations in target scale and distribution, addressing critical limitations of conventional spatial attention approaches. Existing spatial attention methods exhibit suboptimal performance due to their non-discriminative weight allocation across receptive-field regions, as evidenced by overlapping attention map. Such uniform weighting fails to resolve the parameter-sharing constraints inherent to large convolutional kernels. For 3 $\times$ 3 convolutions, RFAConv outperforms ResNet18 and ResNet34 models constructed using other attention mechanisms, a finding supported by the results in Table 3. Unlike convolution operations constructed by other attention mechanisms, RFAConv assigns unique attention weights to each receptive-field slider, enabling the network to highlight key targets, as depicted in Figure 8 for visualization results using the CAM algorithm. Furthermore, enhancing the focus of CA and CBAM on receptive-field spatial features contributes to improved performance. Extending this paradigm, we develop RFCBAM and RFCA by integrating receptive-field spatial feature enhancement into existing attention frameworks, as evidenced in Table 4, where RFCBAMConv and RFCAConv notably improve the accuracy of ResNet18. In comparison to networks built with CBAMConv and CAConv in Table 3, networks constructed by RFCBAMConv and RFCAConv effectively optimize network performance, underscoring the importance of the receptive-field spatial features. Figure 9 further illustrates that RFCBAMConv and RFCAConv emphasize key objects more effectively than CAConv and CBAMConv by adapting different attentional weights for each receptive-field slider.
Experimental validation demonstrates that our proposed method consistently enhances detection performance across benchmark datasets, with quantitative improvements substantiated in Tables 7 and Table 8. On the COCO2017 dataset, the framework achieves significant gains in average precision compared to baseline models, attributed to its capacity to learn location-specific attentional weights for the receptive-field attention method. This design effectively captures spatial variations in target characteristics across complex scenes. However, performance improvements on the VOC dataset remain comparatively limited, likely due to the dataset’s smaller scale and reduced inter-target variability in simpler images. The proposed method learns distinct attentional weights for each receptive-field slider to account for variations in target information at different locations. As a result, the proposed methods demonstrate minimal improvement on VOC data, whereas it significantly enhances network performance on COCO2017.
Empirical observations indicate that while RFAConv offers distinct advantages in semantic segmentation tasks, its performance is not the best. Compared to contemporary attention-enhanced convolutional methods, RFAConv demonstrates measurable improvements over conventional convolution operations, yet its quantitative metrics remain below those of leading attention-based counterparts. This can be explained by the fact that the accuracy of semantic segmentation tasks heavily relies on long-distance information [7], which the design of RFA neglects to capture in spatial terms, leading to its less outstanding performance in semantic segmentation tasks. In contrast, CA performs global pooling in both the H and W directions, thereby partially considering long-distance information and achieving good performance. Although CBAM and CAM also consider long-distance information through global pooling, their compression methods result in significant information loss, leading to suboptimal performance [7]. To further demonstrate the importance of receptive-field spatial features in semantic segmentation tasks, we conduct experiments with RFCBAMConv and RFAConv. The experimental results again demonstrate a significant performance improvement by focusing the attention of existing attention mechanisms on receptive-field spatial features. For example, the comparison between CAConv and RFCAConv highlights this fact.
In summary, the collective experimental findings unequivocally establish RFAConv as a robust solution capable of surpassing standard convolution operations to boost network performance, and only adds a small amount of computational overhead, number of parameters and inference time. Its plug-and-play compatibility enables direct integration into existing architectures through layer-wise substitution of standard convolutions, requiring no structural redesign. Notably, RFAConv assigns unique attention weights to the receptive-field sliders at each position, resulting in significant performance enhancements for datasets with substantial disparities in positional information. Thus, for datasets exhibiting significant discrepancies in location information, such as UAV images, underwater images, and medical images where target characteristics vary across different locations, RFAConv excels in capturing intricate details, thereby enhancing network accuracy.
Limitations. To address the problem of standard convolutional parameter sharing, corresponding attentional weights are learned for each position of the receptive-field slider, which can lead to a larger memory overhead for the network. Therefore devices with limited storage resources may be limited in using our method to construct 3 $\times$ 3 non-shared parameter convolutions. However, it is entirely possible to learn 2 $\times$ 2 attentional weights for each position, which does not completely solve the problem of parameter sharing of standard 3 $\times$ 3 convolution kernels, but can alleviate the problem of parameter sharing of standard convolutions to some extent. This approach improves network performance while maintaining a small increase in memory overhead. Another perfect solution is to build non-square convolutional kernel operations to further flexibly adjust the memory overhead and build non-shared parameter convolutions to boost network performance.
@@ 5 Conclusion
Existing works predominantly concentrate on leveraging spatial attention to improve network performance, exploring its contributions to feature extraction, information selection. Although they indeed show the effectiveness of spatial attention to some degree, they tend to analyze spatial attention from a conventional viewpoint, overlooking the deeper relationship between its intrinsic mechanism and convolutional kernels. This work aims to showcase the efficacy of spatial attention by dissecting the inherent link between spatial attention and the standard convolution from a fresh perspective. It essentially solves the problem of convolution kernel parameter sharing. However, the current performance of spatial attention mechanisms with large convolution kernels remains restricted. Due to the shared attention weights within each receptive-field, the issue of parameter sharing in convolutional kernels is not entirely resolved.
To address this limitation, a novel attention mechanism called Receptive-Field Attention (RFA) is proposed, which emphasizes the receptive-field spatial feature and the significance of each feature within the sliding window. Building on the RFA concept, a Receptive-Field Attention Convolution (RFAConv) is developed to substitute the traditional standard convolution operation. It significantly improves network performance with almost negligible computational overhead and parameter increase, and is expected to replace traditional standard convolutional operations as a core component of neural network architectures for classification, target detection and semantic segmentation tasks. Furthermore, this work underscores the importance of spatial attention mechanisms focusing on receptive-field spatial features to further enhance performance. Therefore, we introduce improved versions of RFCBAM and RFCA based on CBAM and CA, and combine them with standard convolutional operations to design non-shared parameter convolutional operations. Subsequently, we meticulously conduct a large number of classification, object detection, and semantic segmentation experiments on authoritative datasets like ImageNet, COCO, VOC, and Roboflow to thoroughly validate their efficacy and refinement.
Although the proposed method significantly enhances network performance, it requires learning distinct attentional weights for each position of the receptive-field slider to tackle the standard convolutional parameter sharing issue. Consequently, this leads to an increase in the network’s memory overhead. A perfect solution is to build non-square convolutional kernel operations to further flexibly adjust the memory overhead and build non-shared parameter convolutions to boost network performance. Therefore, In the future, we will develop a convolution operation with an arbitrarily shaped convolution kernel to flexibly tune its corresponding memory overhead. Moreover, we anticipate a shift towards increased research on receptive-field spatial features, aiming to leverage the benefits of spatial attention for addressing parameter sharing challenges in standard convolutional operations to dramatically improve network performance.
CRediT authorship contribution statement
Xin Zhang: Conceptualization, Methodology, Validation, Software, Formal analysis, Investigation, Writing-original draft, Writing-review & editing, Visualization. Chen Liu: Resources, Data Curation, Writing-review & editing, Validation. Tingting Song: Methodology, Resources, Writing-review & editing, Project administration, Supervision, Funding acquisition. Degang Yang: Methodology, Resources, Writing-review & editing, Project administration, Supervision, Funding acquisition. Yichen Ye: Data Curation, Resources, Formal analysis, Writing-review & editing. Ke Li: Data Curation, Visualization, Resources, Formal analysis. Yingze Song: Data Curation, Visualization, Resources, Formal analysis.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This research was supported in part by National Natural Science Foundation of China (62441506), in part by Natural Science Foundation of Chongqing (CSTB2022NSCQ-MSX1200), in part by the Science and Technology Research Program of Chongqing Municipal Education Commission (KJQN202200537 and KJZD-M202300502), in part by Chongqing Normal University PhD Start-up Fund (21XLB035), and in part by Chongqing Normal University Postgraduate Research and Innovation Program (YKC24017).